OpenStudy (anonymous):

Suppose i have a function: A*(3-x)+B*A(sqrt(x^2+4)). Find the x-value of the critical point of this cost function, where A,B illustrate the cost. This point will be a function of U

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

the function illustrates the cost function, but how do you find the critical point of it when you have A and B to worry about...i tried differentitiating and i got\[C=-A+BAx(x^2+4)\]

OpenStudy (anonymous):

when C' sorry

OpenStudy (anonymous):

where C is the cost function

OpenStudy (anonymous):

so i let C'=0 and i get \[-A+BAx(x^2+4)=0\]

OpenStudy (anonymous):

\[Bx(x^2+4)=1\]

OpenStudy (anonymous):

Are you sure you differentiated correctly?

OpenStudy (anonymous):

o i forgot the -1/2 aye

OpenStudy (anonymous):

\[C'=-A+\frac{ BAx }{ \sqrt(x^2+4) }\]

OpenStudy (anonymous):

so when C'=0 \[Bx=\sqrt(x^2+4)\]

OpenStudy (anonymous):

Yeah, technically you could solve further.

OpenStudy (anonymous):

\[B^2x^2=x^2+4\]

OpenStudy (anonymous):

\[x^2(B^2-1)-4=0\]

OpenStudy (anonymous):

Ah, difference of squares.

OpenStudy (anonymous):

\[(x-\sqrt(B-1))(x+\sqrt(B-1))=0\]

OpenStudy (anonymous):

\(x>0\)

OpenStudy (anonymous):

so x=sqrt(B-1)?

OpenStudy (anonymous):

ive tried this answer but it says its incorrect

OpenStudy (anonymous):

o wait i see my mistake

OpenStudy (anonymous):

yea i got it!